OpenStudy (ja1):

Below:

OpenStudy (ja1):

Using complete sentences, explain the two ways to simplify the square root of:\[\sqrt{180}\]and provide the simplified form.

OpenStudy (ja1):

I already have the 36 so what is another method?

OpenStudy (skullpatrol):

Where did you get 36?

OpenStudy (anonymous):

9*4

OpenStudy (ja1):

36 * 5 is the greatest perfect square

OpenStudy (anonymous):

180 = 9 * 20 = 9 * 4 * 5

OpenStudy (ja1):

Ok yeah that's the other method

OpenStudy (ja1):

9*20

OpenStudy (ja1):

Thanks all for your two cents :)

OpenStudy (skullpatrol):

$0.02

OpenStudy (ja1):

A receipt will be mailed to you shortly.